if the complimet of an angle is 6 less than 1/3 of its supplement..Find the measure of the angle...
Answer by shree840(260)

let the angle be x
90-x=1/3(180-x)-6
90-x=60-x/3-6
-2x/3=-90+54=-36
x=36*3/2=54
check it'scomplement =36
it's supplement is 180-54=126
1/3=42 42-6=36 Ans